# Easy conditioning far beyond Gaussian

### v3

@@ Appendix B Multivariate unified skew distributions
Multivariate skew normal distributions can be defined in different ways by extending the univariate skew-normal density to the $d$-dimensional case [4]. Here we are mainly interested in the Multivariate Unified Skew Normal distribution, we will explain his stochastic construction and conditioning properties.
@@
then we write $Y\sim SUN_{d,p}(\xi,\gamma,\bar{\omega},\Omega^{*})$ with $\bar{\omega}=\omega 1_{d}$ the vector of diagonal elements of $\omega$, and we say that $Y$ as a multivariate skew unified normal distribution [2].
@@
Let $\mathbf{X}=(\mathbf{X_{1}},\mathbf{X_{2}})^{t}\sim SUN_{d,p}(\xi,\gamma,\bar{\omega},\Omega^{*})$ with $\mathbf{X_{1}}\in\mathbb{R}^{\ell}$ and $\mathbf{X_{2}}\in\mathbb{R}^{m}$ with $\ell+m=d$, $\xi=(\xi_{1},\xi_{2})^{t}$ with $\xi_{1}\in\mathbb{R}^{\ell},~{}\xi_{2}\in\mathbb{R}^{m}$ and $\bar{\omega}_{2}=\text{diag}(\omega_{\ell+1},\ldots,\omega_{d})\in\mathbb{R}^{m\times m}$ The marginal distribution of $\mathbf{X_{2}}$ is
@@ C.2 CKDE Benchmark
The CKDE method used as benchmark, fits a Kernel Density Estimation (KDE) model (using a Gaussian kernel and a bandwidth of 1) on the joint distribution of the response variable $\mathbf{Y}$ and predictor $\mathbf{X}$. It then generates samples from the conditional distribution $\mathbb{P}(\mathbf{Y}|\mathbf{X})$. This is achieved by drawing an oversampled set of joint samples from the KDE model, filtering those to include only the samples close to a given predictor value $\mathbf{X}_{\text{cond}}$ within a specified tolerance (chosen to be 0.1) , and finally selecting the required number of samples (1000) from this filtered set. This process allows obtaining samples from the conditional distribution without needing an analytical form of the joint or conditional distributions.
C.3 Scoring rules
@@ Appendix D Additional figures
Figure 9: Visualization of the joint, marginal and copulas distributions of the Wine data set.
@@
Figure 11: Visualization of the joint, marginal and copulas distributions of the breast cancer data Wisconsin.Red points represent patients diagnosed with malignant, while green points represent patients diagnosed with benign.

### v4

@@ Appendix B Multivariate unified skew distributions
Multivariate skew normal distributions can be defined in different ways by extending the univariate skew-normal density to the $d$-dimensional case ([azzalini1996multivariate]). Here we are mainly interested in the Multivariate Unified Skew Normal distribution, we will explain his stochastic construction and conditioning properties.
@@
then we write $Y\sim SUN_{d,p}(\xi,\gamma,\bar{\omega},\Omega^{*})$ with $\bar{\omega}=\omega 1_{d}$ the vector of diagonal elements of $\omega$, and we say that $Y$ as a multivariate skew unified normal distribution ([arellano2006unification]).
@@
Let $\mathbf{X}=(\mathbf{X_{1}},\mathbf{X_{2}})^{t}\sim SUN_{d,p}(\xi,\gamma,\bar{\omega},\Omega^{*})$ with $\mathbf{X_{1}}\in\mathbb{R}^{\ell}$ and $\mathbf{X_{2}}\in\mathbb{R}^{m}$ with $\ell+m=d$, $\xi=(\xi_{1},\xi_{2})^{t}$ with $\xi_{1}\in\mathbb{R}^{\ell},~\xi_{2}\in\mathbb{R}^{m}$ and $\bar{\omega}_{2}=\text{diag}(\omega_{\ell+1},\ldots,\omega_{d})\in\mathbb{R}^{m\times m}$ . The marginal distribution of $\mathbf{X_{2}}$ is
@@ C.2 CKDE Benchmark
The CKDE method used as benchmark, fits a Kernel Density Estimation (KDE) model (using a Gaussian kernel and a bandwidth of 1 by using the function KernelDensity of the package sklearn in Python) on the joint distribution of the response variables and predictors $\mathbf{X}=(\mathbf{X_{1}},\mathbf{X_{2}})$ . It then generates samples from the conditional distribution $\mathbb{P}(\mathbf{X_{1}}|\mathbf{X_{2}})$. This is achieved by drawing an oversampled set of joint samples from the KDE model, filtering those to include only the samples close to a given predictor value $\mathbf{X}_{\text{cond}}$ within a specified tolerance (chosen to be 0.4 for most of the experiments, except for the experiment on the Wine dataset where it is chosen to be 10), and finally selecting the required number of samples (1000) from this filtered set. This process allows obtaining samples from the conditional distribution without needing an analytical form of the joint or conditional distributions.
C.3 Imputation experiment
The dataset for the imputation experiment on simulated data 5.3 has been generated by sampling $1000$ points from a Gaussian Mixture Copula Model (GMCM) with two components, with means $\mu_{1}=(1,2)^{t}$ and $\mu_{2}=(3,4)^{t}$, weights $w=(0.6,0.4)$ and covariances matrices:
$\Sigma_{1}=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix},~~\Sigma_{2}=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}$
The marginal distributions have been chosen to be from a GMM with respectively means $\mu_{m,1}=(1,2)^{t}$ and $\mu_{2}=(3,4)^{t}$, weights $w_{1}=(0,4,0.6)$ and $w_{2}=(0.4,0.6)$, and standard deviations $\sigma_{1}=(1,1)^{t}$ and $\sigma_{2}=(1,1)^{t}$.
For the experiment with the Breast Cancer Wisconsin data, we focused on a subset of four variables: perimeter standard error (PSE), extreme smoothness (ES), extreme concavity (EC), and extreme concave points (ECP).
In both these experiments, we used the function qmixnorm from the R package KScorrect to compute quantiles of a Gaussian Mixture Model.
The computer used for this experiment is a HP Elite Book 845 with a a AMD Ryzen 7 6800U processor and 16 Go of RAM.
C.4 Scoring rules
@@ Appendix D Additional figures
Figure 9: Visualization of all univariate and bivariate marginal combinations in both original and uniform spaces of the Wine data set.
@@
Figure 11: Visualization of all univariate and bivariate marginal combinations in both original and uniform spaces for Wisconsin breast cancer data. Red points denote patients with malignant tumors, while green points represent those with benign tumors.